Mark and Kelly have the same birthday. Mark is twice as old as Kelly was when Mark was as old as Kelly is now. If Mark is 24 years old now, how old is Kelly?
Answer by ankor@dixie-net.com(22740)

Mark and Kelly have the same birthday.
m = Mark's age
k = Kelly's age
d = (m-k) the difference in their ages
:
Mark is twice as old as Kelly was when Mark was as old as Kelly is now.
m = 2(k - d)
Replace d with (m-k)
m = 2(k -(m-k))
Remove brackets
m = 2(k - m + k)
m = 2(2k - m)
m = 4k - 2m
m + 2m = 4k
3m = 4k
:
If Mark is 24 years old now, how old is Kelly?
3(24) = 4k
72 = 4k
k = 72/4
k = 18 yrs is Kelly's age